OpenStudy (anonymous):

Two groups of students order burritos and tacos at a local restaurant. One order of 3 burritos and 4 tacos costs $11.33. the other order of 9 burritos and 5 tacos costs $23.56. a. Write a system of equations that describes the situation b. Solve by elimination to find the cost of a burrito and the cost of a taco.

OpenStudy (anonymous):

lets say that : the cost of one burrito= x the cost of one taco= y then, 3x+4y=11.33 9x+5y=23.56 that was the system the second question: if u multiply the first equation with -3 it will be -9x-12y=-33.99 9x+5y=23.56 -7y=-10.43 y=1.49 by plugging y in the first equation u will have x=1.79 so the cost of one burrito= 1.79 the cost of one taco= 1.49

OpenStudy (anonymous):

THANK YOU SO MUCH!